# Information theory for data-driven model reduction in physics and biology

### v3

@@ IV Knowing when to stop
For discrete encoding variables $h$, the information bottleneck partitions state space and reduces the dynamics on $x$ to a discrete dynamics on $h$. Such reductions of complex systems to symbolic sequences via partitioning of state space has attracted attention for more than half a century in both theoretical and data-driven contexts [59, 60, 61, 62, 4, 63, 64]. Several works have approached this partition problem from a dynamical systems perspective, linking optimal partitions to eigenfunctions of the (adjoint) transfer operator [33, 65]. In this setting, a central question is “when to stop” [33, 34, 4, 63]: how many states does $h$ need in order to capture statistical properties of the original dynamics?
@@
We consider this question by finding the optimal IB encoder in the limit of low compression, $\beta\gg 1$, but fixed encoding capacity $N_{H}$ (where $H_{t}\in\{0,...,N_{H}-1\}$), i.e. the encoder is only restricted by the number of symbols it can use. An analogous setup was used in the context of renormalization group (RG) transformations in [66, 67, 68], which results in effective model reduction due to the “sloppiness”, or irrelevance, of certain system variables [69, 70]. In this regime, the encoder learned by IB is deterministic; we are learning an optimal hard partition of state space. This can be seen by noting that $I(H_{t};X_{t+\Delta t})=\mathcal{H}(H_{t})-\mathcal{H}(H_{t}|X_{t+\Delta t})$ is maximized when the latter term is zero, which happens when $x_{t}$ unambiguously determines $h_{t}$, i.e. when $p(h_{t}|x_{t})\in\{0,1\}$ for all $x_{t}$. The details of how the encoder is computed are discussed in the next section.
@@
To see this reflected in the information, we again consider an encoding of the initial state into a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. In both the small and large noise scenarios, a variable with $N_{H}=2$ encodes approximately one bit of information (Fig. 3b), corresponding to an $H_{t}$ which distinguishes the two large wells for $x\lessgtr 0$. For large noise this is essentially all the information that can be learned; increasing the capacity of the encoding variable beyond this provides only marginally more information about the future state (Fig. 3c). In the small noise case, the information between the encoding and the future state continues to increase to approximately two bits at $N_{H}=4$, after which it plateaus. The encoding has learned to distinguish each of the four potential wells. These observations are reflected in the transfer operator spectrum shown in Fig. 3d. For small noise, the eigenvalue $\lambda=1$ is nearly four-fold degenerate, indicating the existence of four regions that can evolve independently under $U$, giving rise to four steady state distributions satisfying $U\rho=\rho$. These regions correspond to the potential minima. Hops between the separate minima are exceedingly rare, so that the dynamics essentially take place in the four minima independently. With larger $\sigma$ the degeneracy is lifted, resulting in one dominant subleading eigenvalue followed by a gap. The corresponding eigenfunction is one which is positive (negative) on the right (left) side of the large potential barrier at $x=0$: the only relevant piece of information is which of the large wells the initial condition is contained in, and all other information is lost exponentially quickly.
@@
In many cases, our goal is not only to find a reduced representation $H_{t}$ that is predictive of the future state, but one which may be evolved forward for long times as a surrogate for the full system. This is a practical but much more stringent requirement of our reduced state, and it may result in a different answer to the question of “when to stop.”
@@
In general the dynamics of the reduced variable will be non-Markovian, meaning that its evolution will be determined not only by its current state, but also its past. With a fixed encoder $p(h_{t}|x_{t})$, transition probabilities $p(h_{t+1}|h_{t})$ are induced by decoding $h_{t}$ to $x_{t}$, evolving $x_{t}$ forward in time to $x_{t+1}$, and then encoding $x_{t+1}$ to $h_{t+1}$ (here we assume discrete time dynamics with $\Delta t=1$ for simplicity). Information theory provides a natural quantification of the (non-)Markovianity of the dynamics as measured by the conditional mutual information $I(H_{t-1},H_{t+1}|H_{t})$ (see SI Section 3 and alternative approaches in Refs. [71, 72, 63]). This quantity tells us how much additional information the past state $H_{t-1}$ contains about the future $H_{t+1}$, on top of what is contained in the present state $H_{t}$.
@@
The role of memory is illustrated in the two cases shown in Fig. 4b. In both cases, the current state is in the right well, so that $h_{t}=1$. However, more precise information about the state (and hence its future) is revealed by knowing the system’s past. If $h_{t-1}=0$ (so the state was in the left well), it is likely that the full state is near the boundary between wells, so that $h_{t+1}=0$ is more likely than if $h_{t-1}=1$, which would suggest that the state is deep in the right well. This behavior can be quantified by computing the transfer probabilities $p(h_{t+1}|h_{t},h_{t-1})$ (Fig. 4c). These differ significantly from $p(h_{t+1}|h_{t})$, which would not be the case for Markovian dynamics. From these, we see that the primary contribution to the non-Markovian dynamics occurs when $(h_{t-1},h_{t})=(0,1)$ or $(1,0)$, which can only happen when the state is very near to the boundary between the wells. As $N_{H}$ increases, the likelihood that the state is near the boundary of two regions with different $h_{t}$ labels will increase, which will increase the non-Markovianity of the system (Fig. 4d-e). However, at the same time the different $h_{t}$ regions will shrink in size so that $h_{t-1}$ will no longer provide any meaningful information about the system’s location $x_{t}$. The trade-off between these two effects leads to the observed non-monotonic behavior of $I(H_{t-1},H_{t+1}|H_{t})$.
V Data-driven discovery of slow variables
These variables coincide with left eigenfunctions of the transfer operator, as discussed in Section II. We verify this theoretical result numerically in cases where the IB objective (3) can be solved exactly by using an iterative scheme known as the Blahut-Arimoto algorithm [38, 58] (see SI Section 3). By taking the logarithm of the encoder as it begins to deviate from the trivial uniform encoder above $\beta_{1}$ we can confirm that it depends on $x$ only through $\phi_{1}(x)$ (SI Fig. S3). This can be independently verified by studying the stability of the trivial encoder with respect to perturbations in the parameters $f_{n}(h_{t})$ in Eq. 7, where we see that the encoder becomes unstable to perturbations in $f_{1}$, as predicted by our theoretical results (SI Section 4).
The utility of exact IB for variable discovery is limited because it requires knowledge of the exact conditional distribution $p(x_{t+\Delta t}|x_{t})$ which is difficult to estimate in many real-world scenarios. Fortunately however, the IB optimization problem can be replaced by an approximate variational objective introduced in Ref. [41] that can be solved with neural networks and which we refer to as variational IB. This is achieved by introducing a tractable Ansatz for the form of $p(h_{t}|x_{t})$; we show in SI Section 4 that this does not change the encoder’s dependence on transfer operator eigenfunctions.
Figure 5: Variational IB for high-dimensional simulated fluid flow. (a) Fluid flow past a disk with uniform in-flow velocity $v_{0}$ exhibits vortex shedding behind the object in a so-called von Kármán street. The state of the system is given by a spatially varying two-component vector field $\mathbf{v}(\mathbf{x})$. (b) Dynamics in latent space (blue) traverse a nearly circular trajectory (see Supplementary Movie 1). For comparison we show the evolution of the mode amplitudes obtained by projecting the velocity field onto the first DMD mode (green). (c) Comparison of the first Koopman mode obtained from DMD ($\mathbf{m}^{(1)}$) and from VIB ($\mathbf{m}^{(\text{IB})}$). Koopman modes from VIB are computed as gradients of the latent encoding variables as described in the main text. Red corresponds to positive values and blue to negative; the magnitudes of the modes are not directly comparable. (d) Intensity field $I(x)$ from a video of water flowing past an obstacle; video from Ref. [73]. (e) Trajectory of the system in latent space. (f) Gradients of the latent encodings with respect to the input field $I(x)$. (g-i) Same as (d-f), but using satellite images of clouds near the island of Guadalupe as our state $I(x)$.
Our core result linking the behavior of the learned encoder with transfer operator eigenfunctions remains valid even for variational IB applied to high-dimensional systems, which we show by considering a simulated data of fluid flow past a disk [74]. The state of the system is given by a two-dimensional velocity field $\mathbf{v}(\mathbf{x})\in\mathbb{R}^{2\times N_{\text{pixels}}}$, where $N_{\text{pixels}}\sim\mathcal{O}(10^{5})$ (Fig. 5a). Fluid flows in from the left boundary with a constant velocity $v_{0}\hat{e}_{x}$ past a disk of unit diameter. At Reynolds number $\text{Re}\gtrsim 150$, the fluid undergoes periodic vortex shedding behind the disk, forming what is known as a von Kármán street.
Our learned latent variables are oscillatory with the correct frequency as shown in Fig. 5b. However, this alone does not indicate whether we are learning the correct eigenfunctions. Because the system is well approximated by linear dynamics, eigenfunctions of the adjoint transfer operator are linear functions of the state variable, and may hence be represented as $\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle$, where angled brackets denote an average over space and $\mathbf{m}^{(n)}(\mathbf{x})$ are so-called Koopman modes. The true $\mathbf{m}^{(n)}$ (and hence the eigenfunctions) can be thus computed via dynamic mode decomposition (DMD) [15, 16] as described in the SI. For the functions $h[\mathbf{v}]$ learned using variational IB, we can infer the corresponding $\mathbf{m}^{(\text{IB})}$ using gradients of the neural network with respect to the input field. Fig. 5c shows these inferred modes $\mathbf{m}^{(\text{IB})}$ compared to the true mode $\mathbf{m}^{(1)}$ (see SI Section 5 for details). This shows that variational IB not only recovers the essential oscillatory nature of the dynamics, but does so by learning the correct slowly varying functions of the state variable given by the adjoint transfer operator eigenfunctions.
We also apply variational IB to a von Kármán street arising due to air flow around Guadalupe Island, which was imaged by a National Oceanic and Atmospheric Administration (NOAA) satellite [75] (Fig. 5g). Although the video consists of only 62 frames and less than one full oscillation, the variational IB neural network learns latent variables which capture this oscillation and have the expected dependence on the input variables (Fig. 5h-i). As in the first experimental example, the gradients of the encoding variables show the glide symmetry of $m_{x}$ due to the symmetry of the intensity pattern in Fig. 5g.
@@ V Data-driven discovery of slow variables
Figure 6: IB as part of an equation learning pipeline. (a) Kymographs of the sine-Gordon system, discussed in the text. Left shows the field $\phi(x)$ while the right shows time-derivative of the field. Initial conditions are shown above as curves. (b) Information content contained in the encoding for varying dimensionality of the embedding space. While the information appears to plateau at $D^{*}=2$, this does not necessarily mean a “proper” embedding dimension has been reached. (c) Estimated fractal dimension of the embedded dynamics for varying embedding dimension. The full fractal dimension of $\approx 2$ is only attained with an encoding dimension $D^{*}=3$. (d) Error of the predictions for the latent variable evolution produced by the dynamical system learned by SINDy. The error reaches a minimum at $D^{*}=4$. Interestingly the error begins to rise for $d>4$, likely due to the fact that the IB encoding begins to store other features of the full state variable which are irrelevant for the dynamics. (e) For $d=2$, the encoding manages to project the full torus into two dimensions by cutting and unfolding it (left). This introduces discontinuous jumps in the evolution of $h$. (f) Time evolution of both components of $h$, with highlighted part corresponding to the trajectory highlighted in (e). (g) Trajectory for a four-dimensional latent variable, visualized by projected onto the first three principal components. True trajectory shown in black, trajectory predicted by SINDy shown in blue with the equations given in the main text. (h) Predicted trajectory for all the degrees of freedom of $h$.
VI Learning the dynamics of latent variables
In Section I we showed how using an information-theoretic measure of predictability reduces the problem of model reduction to the problem of identifying relevant variables. In this setting, the dynamical rule governing the evolution of $H_{t}$ is not explicitly learned, but is instead induced by the choice of encoding variable. We now show how variational IB can be incorporated into a complete model reduction pipeline by combining it with sparse equation learning methods such as SINDy [76, 77]. We consider a system described by an underdamped, driven, one-dimensional PDE known as the sine-Gordon equation [78] (Fig. 6a; see SI Section 9). For a particular choice of driving frequency and strength, this system undergoes quasiperiodic dynamics, and hence lives on a toroidal manifold.
@@ VI Learning the dynamics of latent variables
The “fraying” evident at the boundary of the torus embedded in two dimensions leads to a fractal dimension which is less than two, although this difference may disappear in the limit of infinite data. At three dimensions, the latent manifold is two dimensions everywhere (Fig. 6c). However, while the torus may be embedded in three dimensions, it does not admit simple, linear dynamics until $\dim H_{t}=4$. In this dimension, SINDy is able to identify the correct linear equations of motion
@@
VII Relevant variable discovery in cyanobacterial populations

### v4

@@ IV.1 Knowing when to stop
For discrete encoding variables $h$, the information bottleneck partitions state space and reduces the dynamics on $x$ to a discrete dynamics on $h$. Such reductions of complex systems to symbolic sequences via partitioning of state space have attracted attention for more than half a century in both theoretical and data-driven contexts [83, 84, 85, 86, 87, 88, 89, 90]. Several works have approached this partition problem from a dynamical systems perspective, linking optimal partitions to eigenfunctions of the (adjoint) transfer operator [56, 57]. In this setting, a central question is “when to stop” [56, 58, 87, 88]: how many states does $h$ need in order to capture statistical properties of the original dynamics?
@@
We consider this question by finding the optimal IB encoder in the limit of low compression, $\beta\gg 1$, but fixed encoding capacity $N_{H}$ (where $H_{t}\in\{0,...,N_{H}-1\}$), i.e. the encoder is only restricted by the number of symbols it can use. An analogous setup was used in the context of renormalization group (RG) transformations in [91, 92, 93], which results in effective model reduction due to the “sloppiness”, or irrelevance, of certain system variables [94, 95]. In this regime, the encoder learned by IB is deterministic; we are learning an optimal hard partition of state space. This can be seen by noting that $I(H_{t};X_{t+\Delta t})=\mathcal{H}(H_{t})-\mathcal{H}(H_{t}|X_{t+\Delta t})$ is maximized when the latter term is zero, which happens when $x_{t}$ unambiguously determines $h_{t}$, i.e. when $p(h_{t}|x_{t})\in\{0,1\}$ for all $x_{t}$. The optimal encoder is computed using the Blahut-Arimoto algorithm (Ref. [41, 82] and SI Section S1).
@@
To see this reflected in the information, we again consider an encoding of the initial state into a discrete variable $H_{t}\in\{0,...,N_{H}-1\}$. In both the small and large noise scenarios, a variable with $N_{H}=2$ encodes approximately one bit of information (Fig. 3b), corresponding to an $H_{t}$ which distinguishes the two large wells for $x\lessgtr 0$. For large noise this is essentially all the information that can be learned; increasing the capacity of the encoding variable beyond this provides only marginally more information about the future state (Fig. 3c). In the small noise case, the information between the encoding and the future state continues to increase to approximately two bits at $N_{H}=4$, after which it plateaus. The encoding has learned to distinguish each of the four potential wells. These observations are reflected in the transfer operator spectrum shown in Fig. 3d. For small noise, the eigenvalue $\lambda=1$ is nearly four-fold degenerate, indicating the existence of four regions that can evolve independently under $U$, giving rise to four steady state distributions satisfying $U\rho\simeq\rho$. These regions correspond to the potential minima. Hops between the separate minima are exceedingly rare, so that the dynamics essentially take place in the four minima independently. With larger $\sigma$ the degeneracy is lifted, resulting in one dominant subleading eigenvalue followed by a gap. The corresponding eigenfunction is one which is positive (negative) on the right (left) side of the large potential barrier at $x=0$: the only relevant piece of information is which of the large wells the initial condition is contained in, and all other information is lost exponentially quickly.
@@
In many cases, our goal is not only to find a reduced representation $H_{t}$ that is predictive of the future state, but one which may be evolved forward for long times as a surrogate for the full system. This is a practical but much more stringent requirement of our reduced state, and it may result in a different answer to the question of “when to stop” which we explore in the following sections.
IV.2 Quantifying non-Markovianity
This section is dedicated to the Markovianity of the latent variables (or the lack thereof), assuming an underlying Markovian dynamics. In practice, this assumption may not hold, for instance when experimental datasets only contain a partial description of the systems. In Sec. V.2, we discuss this situation as well as potential remedial strategies.
@@ IV.2 Quantifying non-Markovianity
In general the dynamics of the reduced variable will be non-Markovian, meaning that its evolution will be determined not only by its current state, but also its past. With a fixed encoder $p(h_{t}|x_{t})$, transition probabilities $p(h_{t+1}|h_{t})$ are induced by decoding $h_{t}$ to $x_{t}$, evolving $x_{t}$ forward in time to $x_{t+1}$, and then encoding $x_{t+1}$ to $h_{t+1}$ (here we assume discrete time dynamics with $\Delta t=1$ for simplicity). Information theory provides a natural quantification of the (non-)Markovianity of the dynamics as measured by the conditional mutual information (SI Section S3):
$\text{NM}(H_{t})\equiv I(H_{t-1},H_{t+1}|H_{t}).$ (20)
This quantity tells us how much additional information the past state $H_{t-1}$ contains about the future $H_{t+1}$, on top of what is contained in the present state $H_{t}$ (see alternative approaches in Refs. [96, 97, 88]).
Figure 4: Learned collective variables are generically non-Markovian (a) Conditional mutual information $I(H_{t-\Delta t},H_{t+\Delta t}|H_{t})$ as a function of $N_{H}$. (b) Two possible scenarios (left and right) with the same value of $h_{t}$, but different $h_{t-1}$. Here, the circles represent the value of the full state $x_{t}$ while the colored regions represent the value of the reduced state $h_{t}\in\{1,2\}$ which are mapped to for each $x_{t}$. (c) Illustration of the transition probability which takes into account the previous state $h_{t-1}$ (left), compared to the Markovian transition probability obtained by marginalizing out $h_{t-1}$ (right). Black means the transition is likely, white means it is unlikely. In a Markovian system, these transition probabilities would coincide. (d-e) Same as in (h), but for $N_{H}=4$ states.
@@
The role of memory is illustrated in the two cases shown in Fig. 4b. In both cases, the current state is in the right well, so that $h_{t}=1$. However, more precise information about the state (and hence its future) is revealed by knowing the system’s past. If $h_{t-1}=0$ (so the state was in the left well), it is likely that the full state is near the boundary between wells, so that $h_{t+1}=0$ is very likely. However, $h_{t+1}=0$ is very unlikely when $h_{t-1}=1$, which happens when the state is deep in the right well. This behavior can be quantified by computing the transfer probabilities $p(h_{t+1}|h_{t},h_{t-1})$ (Fig. 4c). These differ significantly from $p(h_{t+1}|h_{t})$, which would not be the case for Markovian dynamics. The primary contribution to the non-Markovian dynamics occurs when $(h_{t-1},h_{t})=(0,1)$ or $(1,0)$, which occurs when the state is near the boundary between the wells. As $N_{H}$ increases, the likelihood that the state is near the boundary of two regions with different $h_{t}$ labels will increase, which will increase the non-Markovianity of the system (Fig. 4d-e). However, at the same time the different $h_{t}$ regions will shrink in size so that $h_{t-1}$ will no longer provide any meaningful information about the system’s location $x_{t}$. The trade-off between these two effects leads to the observed non-monotonic behavior of $I(H_{t-1},H_{t+1}|H_{t})$.
IV.3 Learning effective dynamics of latent variables
In practice, latent variable identification is often only the first step of a procedure that finds the variables and their dynamics (Section II). We now illustrate how our framework forms the basis for a complete model discovery pipeline by using sparse equation learning to extract relevant variable dynamics [43, 98]. To learn continuous latent variables, we solve the IB optimization problem using an approximate variational objective introduced in Ref. [47] that can be solved with neural networks and which we refer to as variational IB. This approach is discussed in more detail in the following Section V and in SI Section S5.
@@ IV.3 Learning effective dynamics of latent variables
Figure 5: “When to stop” depends on what you want. (a) Kymographs of the sine-Gordon system, discussed in the text. Left shows the field $\phi(x)$ while the right shows time-derivative of the field. Initial conditions are shown above as curves. (b) Information content contained in the encoding for varying dimensionality of the embedding space. While the information appears to plateau at $D^{*}=2$, this does not necessarily mean a “proper” embedding dimension has been reached. (c) Estimated fractal dimension of the embedded dynamics for varying embedding dimension. The full fractal dimension of $\approx 2$ is only attained with an encoding dimension $D^{*}=3$. (d) Error of the predictions for the latent variable evolution produced by the dynamical system learned by SINDy. The error reaches a minimum at $D^{*}=4$. (e) For $d=2$, the encoding manages to project the full torus into two dimensions by cutting and unfolding it (left). This introduces discontinuous jumps in the evolution of $h$. (f) Time evolution of both components of $h$, with highlighted part corresponding to the trajectory highlighted in (e). (g) Trajectory for a four-dimensional latent variable, visualized by projected onto the first three principal components. True trajectory shown in black, trajectory predicted by SINDy shown in blue with the equations given in the main text. (h) Predicted trajectory for all the degrees of freedom of $h$.
We consider a system described by an underdamped, driven, one-dimensional PDE known as the sine-Gordon equation [99] (Fig. 5a; see SI Section S9):
$\partial_{t}^{2}\phi-\partial_{x}^{2}\phi+\sin\phi=\varepsilon[-\alpha\partial_{t}\phi+\Gamma\cos(\omega t)].$ (21)
For a particular choice of driving frequency $\omega$ and strength $\varepsilon\Gamma$, this system undergoes quasiperiodic dynamics, and hence lives on a toroidal manifold.
IV.3.1 When to stop?
@@
The “fraying” visible at the boundary of the torus embedded in two dimensions leads to a fractal dimension which is less than two, although this difference may disappear in the limit of infinite data. At three dimensions, the latent manifold is two dimensions everywhere (Fig. 5c). However, while the torus may be embedded in three dimensions, it does not admit simple, linear dynamics until $\dim H_{t}=4$.
IV.3.2 Symbolic equation learning
In dimension $\dim H_{t}=4$, SINDy is able to identify the correct linear equations of motion
@@ IV.3.2 Symbolic equation learning
V Applications
V.1 von Kármán streets: variable discovery from YouTube videos
These variables coincide with left eigenfunctions of the transfer operator, as discussed in Section III. We verify this theoretical result numerically in cases where the IB objective Eq. (11) can be solved exactly by using an iterative scheme known as the Blahut-Arimoto algorithm [41, 82] (SI Section S1 and S4). By taking the logarithm of the encoder as it begins to deviate from the trivial uniform encoder above $\beta_{1}$ we can confirm that it depends on $x$ only through $\phi_{1}(x)$ (SI Fig. S5a-d). This can be independently verified by studying the stability of the trivial encoder with respect to perturbations in the parameters $f_{n}(h_{t})$ in Eq. (16), where we see that the encoder becomes unstable to perturbations in $f_{1}$, as predicted by our theoretical results (SI Fig. S5e-i).
The utility of exact IB for variable discovery is limited because it requires knowledge of the exact conditional distribution $p(x_{t+\Delta t}|x_{t})$ which is difficult to estimate in many real-world scenarios. Alternatively, as in the previous section, variational IB may be used to solve the IB objective function directly from data. This is achieved by introducing a tractable Ansatz for the form of $p(h_{t}|x_{t})$; we show in SI Section S4 that this does not change the encoder’s dependence on transfer operator eigenfunctions.
Figure 6: Variational IB neural networks learn transfer-operator eigenfunctions in high-dimensional fluid flows. (a) Fluid flow past a disk with uniform in-flow velocity $v_{0}$ exhibits vortex shedding behind the object in a so-called von Kármán street. The state of the system is given by a spatially varying two-component vector field $\mathbf{v}(\mathbf{x})$. (b) Dynamics in latent space (blue) traverse a nearly circular trajectory (see Supplementary Movie 1). For comparison we show the evolution of the mode amplitudes obtained by projecting the velocity field onto the first DMD mode (green). (c) Comparison of the first Koopman mode obtained from DMD ($\mathbf{m}^{(1)}$) and from VIB ($\mathbf{m}^{(\text{IB})}$). Koopman modes from VIB are computed as gradients of the latent encoding variables as described in the main text. Red corresponds to positive values and blue to negative; the magnitudes of the modes are not directly comparable. (d) Intensity field $I(x)$ from a video of water flowing past an obstacle; video from Ref. [100]. (e) Trajectory of the system in latent space. (f) Gradients of the latent encodings with respect to the input field $I(x)$. (g-i) Same as (d-f), but using satellite images of clouds near the island of Guadalupe as our state $I(x)$.
Our core result linking the behavior of the learned encoder with transfer operator eigenfunctions remains valid even for variational IB applied to high-dimensional systems, which we show by considering a simulated data of fluid flow past a disk [101]. The state of the system is given by a two-dimensional velocity field $\mathbf{v}(\mathbf{x})\in\mathbb{R}^{2\times N_{\text{pixels}}}$, where $N_{\text{pixels}}\simeq\mathcal{O}(10^{5})$ (Fig. 6a). Fluid flows in from the left boundary with a constant velocity $v_{0}\hat{e}_{x}$ past a disk of unit diameter. At Reynolds number $\text{Re}\gtrsim 150$, the fluid undergoes periodic vortex shedding behind the disk, forming what is known as a von Kármán street.
Our learned latent variables are oscillatory with the correct frequency as shown in Fig. 6b. However, this alone does not indicate whether we are learning the correct eigenfunctions. Because the system is well approximated by linear dynamics, eigenfunctions of the adjoint transfer operator are linear functions of the state variable, and may hence be represented as $\phi_{n}[\mathbf{v}]=\langle\mathbf{v}(\mathbf{x}),\mathbf{m}^{(n)}(\mathbf{x})\rangle$, where angled brackets denote an average over space and $\mathbf{m}^{(n)}(\mathbf{x})$ are so-called Koopman modes. The true $\mathbf{m}^{(n)}$ (and hence the eigenfunctions) can be thus computed via dynamic mode decomposition (DMD) [21, 22] as described in the SI. For the functions $h[\mathbf{v}]$ learned using variational IB, we can infer the corresponding $\mathbf{m}^{(\text{IB})}$ using gradients of the neural network with respect to the input field. Fig. 6c shows these inferred modes $\mathbf{m}^{(\text{IB})}$ compared to the true mode $\mathbf{m}^{(1)}$ (SI Section S6). This shows that variational IB not only recovers the essential oscillatory nature of the dynamics, but does so by learning the correct slowly varying functions of the state variable given by the adjoint transfer operator eigenfunctions.
Finally, we apply variational IB to a von Kármán street arising due to air flow around Guadalupe Island, which was imaged by a National Oceanic and Atmospheric Administration (NOAA) satellite [102] (Fig. 6g). Although the video consists of only 62 frames and less than one full oscillation, the variational IB neural network learns latent variables which capture this oscillation and have the expected dependence on the input variables (Fig. 6h-i). In general the performance of VIB in the low-data limit is susceptible to overfitting [103]; here this effect is mitigated by the low effective dimensionality of the system. As in the first experimental example, the gradients of the encoding variables show the glide symmetry of $m_{x}$ due to the symmetry of the intensity pattern in Fig. 6g.
V.2 Cyanobacteria: latent variable discovery in experimental data